OpenStudy (anonymous):

How do I simplify: (2x+3/x^2-1) + (3x-4/x^2-4)?

OpenStudy (anonymous):

\[\frac{2(x+1) + 1}{(x-1)(x+1)} + \frac{3(x-2) +2}{(x-2)(x+2)}\] \[\frac{2}{x-1} +\frac{1}{(x+1)(x-1)} + \frac{3}{x+2} +\frac{2}{(x-2)(x+2)} \]

OpenStudy (anonymous):

The answer is 5x-1/x^2-1

OpenStudy (anonymous):

...